# Optimizing Integrated Terrestrial and Non-Terrestrial Networks Performance with Traffic-Aware Resource Management

### v1

@@ V-E Analysis on the Network Sum Log-Throughput
In this section, we analyse the SLT achieved by the algorithms under investigation throughout the day and, in particular, how they adapt the network resources to the daily variations of the traffic load. Figure 8 shows the daily profile of the SLT achieved by the schemes under investigation, and Figure 9 presents the relative SLT gain for BLASTER, HEURISTIC and 3GPP-NTN compared to 3GPP-TN. Remember than in 3GPP-TN there are no satellites available to serve UEs perceiving low RSRP from terrestrial MBSs. In addition,, the total bandwidth available for the terrestrial network is only $10$ MHz, which leads to low UE throughput, especially during busy hours. As expected, we observe an SLT performance improvement with the integration of satellites into the network. Indeed, LEO satellites are able to provide service to cell-edge UEs, which do not perceive a signal strong enough to be served from the terrestrial tier, thereby increasing the network SLT.
@@
In more detail, in Figure 8, we can see that both BLASTER and the proposed HEURISTIC outperform the 3GPP-NTN benchmark, with an average SLT increase of approximately $6\,\%$ across the day, highlighted in Figure 9. Moreover, and although the SLT gain for 3GPP-NTN in Figure 8 is more visually discernible during high-traffic hours, note that its relative gain with respect to 3GPP-TN, illustrated in Figure 9 does not change drastically and remains around $2\%$ throughout the day.
@@
In low traffic, Figure 9 underlines that BLASTER outperforms HEURISTIC in terms of SLT. This is due to the different association methods used for both proposed frameworks. As detailed in Section V-D, the association criterion to the satellite in HEURISTIC is more lenient, which results to more UEs sharing the resources of the satellite, leading to a deteriorated throughput for its UEs and a worse SLT performance. Note that the improvement compared to the 3GPP benchmarks is more apparent during high traffic hours, typically midday to end of evening than low traffic hours. Indeed, as detailed before, the bonus of both the HEURISTIC and BLASTER during low traffic is on reducing energy consumption, which explains the mitigated improvement of the network SLT.
@@
During high traffic, both algorithms strive to balance the traffic load to maximize the SLT, which explains the striking improvement with respect to 3GPP-NTN and 3GPP-TN. That gain is due to various reasons. First of all, the UE-MBS association methods of HEURISTIC and BLASTER are designed to ensure a proportional fair resource allocation, which increases the SLT. In addition, the dynamic split of the network bandwidth based on the time-varying fraction of UEs associated to the satellite network (see Proposition 2) allows the proposed HEURISTIC and BLASTER to astutely distribute the frequency resources, differently than the 3GPP benchmarks [29]. In fact, since the majority of UEs are associated to a terrestrial MBS (see Figure 7), a larger share of the resources are allocated to the terrestrial tier, which leads to a larger provided data-rate than the one achievable following the 3GPP benchmarks. By splitting the bandwidth based on the fraction of UEs associated to the satellites, we observe an enhancement of the mean SLT by at least $8\%$ and $6\%$ for both the proposed frameworks compared to 3GPP-TN and 3GPP-NTN respectively.
@@
V-F Analysis on the Terrestrial Network Energy Consumption
In this section, we study the performance of the proposed BLASTER and HEURISTIC in terms of energy usage. To this end, Figure 10 displays the terrestrial network energy consumption throughout the day for the various algorithms under investigation. The red dotted line represents the energy consumption level for both 3GPP-NTN and 3GPP-TN settings, for which the terrestrial network energy consumption is at maximum level through the entire day, as they do not integrate neither power control nor MBS shutdown. As discussed in Section V-D, with BLASTER and HEURISTIC, the satellite network serves a larger proportion of UEs during low traffic than the standard 3GPP-NTN, which facilitates the shutdown of terrestrial MBSs. This is apparent in Figure 10, as, in low traffic, the terrestrial network energy consumption sees an average decrease of approximately $67\,\%$ and $62\,\%$ for the BLASTER and HEURISTIC settings with respect to the benchmark 3GPP-NTN and 3GPP-TN, respectively. We also notice that even though, during the low-traffic hours, the HEURISTIC is characterized by a larger share of UEs associated to the satellite network than the BLASTER, it still leads to a larger terrestrial network energy consumption. Interestingly, the transmit power optimization of BLASTER allows for a greater reduction of the transmit power of the terrestrial network MBSs. Indeed, the transmit power vector update step for BLASTER (21) - (26) is specifically designed to solve the problem, reducing the terrestrial MBSs transmit power efficiently while the HEURISTIC updates the transmit power for terrestrial MBSs based on a rule of thumb, which suffices to reduce the network energy consumption. This also explains the striking difference in the high-traffic scenario, as the average energy consumed by the network is reduced by around $53\,\%$ and $14\,\%$ for BLASTER and HEURISTIC respectively, compared to the terrestrial network energy consumption in the 3GPP benchmarks. This result further underlines the flexible nature of both proposed frameworks, adjusting their behaviours to the traffic state and its specific demand.
Figure 10: Daily profile of the terrestrial network energy consumption.
@@ VI Conclusions
[…]
Proof of Proposition 1. First, we inject (12) into formula (13) to get:
@@
Indeed, developing the left component of equation (30), we get:
@@
Then, focusing on the right component of the equation, we have:
@@
$=\begin{bmatrix}\mu_{1}\left(\sum\limits_{j=1}^{L}x^{*}_{1j}\beta_{1j}p_{j}\right)&\cdots&\cdots\\ \vdots&\ddots&\vdots\\ \cdots&\cdots&\mu_{K}\left(\sum\limits_{j=1}^{L}x^{*}_{Kj}\beta_{Kj}p_{j}\right)\\ \end{bmatrix}$

### v2

@@ V-E Analysis on the Network Sum Log-Throughput
In this section, we analyse the SLT achieved by the algorithms under investigation throughout the day and, in particular, how they adapt the network resources to the daily variations of the traffic load. Fig. 9 shows the daily profile of the SLT achieved by the schemes under investigation, and Fig. 10 presents the relative SLT gain for BLASTER, HEURISTIC and 3GPP-NTN compared to 3GPP-TN. Remember that in 3GPP-TN there are no satellites available to serve UEs perceiving low RSRP from terrestrial MBSs. In addition, the total bandwidth available for the TN is only $10$ MHz, which leads to low UE throughput, especially during busy hours. As expected, we observe an SLT performance improvement with the integration of satellites into the network. Indeed, LEO satellites are able to provide service to cell-edge UEs, which do not perceive a signal strong enough to be served from the terrestrial tier, thereby increasing the network SLT.
@@
In more detail, in Fig. 9, we can see that both BLASTER and the proposed HEURISTIC outperform the 3GPP-NTN benchmark, with an average SLT increase of approximately $6\,\%$ across the day, highlighted in Fig. 10. Moreover, and although the SLT gain for 3GPP-NTN in Fig. 9 is more visually discernible during high-traffic hours, note that its relative gain with respect to 3GPP-TN, illustrated in Fig. 10 does not change drastically and remains around $2\%$ throughout the day.
@@
In low traffic, Fig. 10 underlines that BLASTER outperforms HEURISTIC in terms of SLT. This is due to the different association methods used for both proposed frameworks. As detailed in Section V-D, the association criterion to the satellite in HEURISTIC is more lenient, which results to more UEs sharing the resources of the satellite, leading to a deteriorated throughput for its UEs and a worse SLT performance. Note that the improvement compared to the 3GPP benchmarks is more apparent during high traffic hours, typically midday to end of evening than low traffic hours. Indeed, as detailed before, the onus of both the HEURISTIC and BLASTER during low traffic is on reducing energy consumption, which explains the mitigated improvement of the network SLT.
@@
During high traffic, both algorithms strive to balance the traffic load to maximize the SLT, which explains the striking improvement with respect to 3GPP-NTN and 3GPP-TN. That gain is due to various reasons. First of all, the UE-MBS association methods of HEURISTIC and BLASTER are designed to ensure a proportional fair resource allocation, which increases the SLT. In addition, the dynamic split of the network bandwidth based on the time-varying fraction of UEs associated to the satellite network (see Proposition 2) allows the proposed HEURISTIC and BLASTER to astutely distribute the frequency resources, differently than the 3GPP benchmarks [40]. In fact, since the majority of UEs are associated to a terrestrial MBS (see Fig. 8), a larger share of the resources are allocated to the terrestrial tier, which leads to a larger provided data-rate than the one achievable following the 3GPP benchmarks. By splitting the bandwidth based on the fraction of UEs associated to the satellites, we observe an enhancement of the mean SLT by at least $8\%$ and $6\%$ for both the proposed frameworks compared to 3GPP-TN and 3GPP-NTN respectively.
@@
V-F Analysis on the Network Energy Consumption
In this section, we study the performance of the proposed BLASTER and HEURISTIC in terms of energy usage. To this end, Fig. 11 displays the TN energy consumption throughout the day for the various algorithms under investigation. The red dotted line represents the energy consumption level for both 3GPP-TN and 3GPP-NTN settings, for which the TN energy consumption is at maximum level through the entire day, as they do not integrate neither power control nor MBS shutdown. Also, the black dotted line represents the same setting as 3GPP-TN, but with the added mechanism of shutting down all inactive MBSs when they have no UEs, denoted 3GPP--ENERGY SAVING. This setting reduces the average daily energy consumption by $17$ compared to 3GPP benchmarks, due to its ability to shut down MBSs, which leads to a more efficient energy use. As discussed in Section V-D, with BLASTER and HEURISTIC, the satellite network serves a larger proportion of UEs during low traffic than the standard 3GPP-NTN, which facilitates the shutdown of terrestrial MBSs. This is apparent in Fig. 11, as, in low traffic, the TN energy consumption sees an average decrease of approximately $67\,\%$ and $54\,\%$ for the BLASTER and HEURISTIC settings with respect to the benchmark 3GPP-NTN and 3GPP-TN, respectively. In comparison, for the 3GPP--ENERGY SAVING configuration, BLASTER and HEURISTIC experience a decrease of approximately $49\,\%$ and $31\,\%$, respectively. We also notice that, even though, during the low-traffic hours, the HEURISTIC is characterized by a larger share of UEs associated to the satellite network than the BLASTER, it still leads to a larger TN energy consumption. Interestingly, the transmit power optimization of BLASTER allows for a greater reduction of the transmit power of the TN MBSs. Indeed, the transmit power vector update step for BLASTER (21) - (26) is specifically designed to solve the optimization problem, reducing the terrestrial MBSs transmit power efficiently while the HEURISTIC updates the transmit power for terrestrial MBSs based on a rule of thumb, which suffices to reduce the network energy consumption. This also explains the striking difference in the high-traffic scenario, as the average energy consumed by the network is reduced by around $56\,\%$ and $14\,\%$ for BLASTER and HEURISTIC respectively, compared to the TN energy consumption in the 3GPP benchmarks. This result further underlines the flexible nature of both proposed frameworks, adjusting their behaviours to the traffic state and its specific demand. Moreover, Fig. 12 depicts the energy consumption of the satellite throughout the day for the settings detailed above. We notice that BLASTER and HEURISTIC considerably reduce the daily average energy consumption of the satellite by up to $80\%$ and $90\%$ respectively. It is important to note that the 3GPP benchmark is configured such that it utilizes the full $30$ MHz of bandwidth made available throughout the day, which leads to consistently higher energy consumption. In contrast, BLASTER and HEURISTIC use a smaller portion of the total bandwidth, as highlighted in Fig. 8 which results in lower energy usage. Additionally, by comparing Fig. 11 and Fig. 12, we notice that the energy consumption of a single satellite is negligible compared to the TN (less than $0.1\%$ of the total terrestrial energy use), which further validates our earlier assumption in Section III that the satellite energy consumption is insignificant in comparison to the TN.
Figure 11: Daily profile of the TN energy consumption.
Figure 12: Daily profile of the satellite energy consumption
@@ VI Conclusions
In this paper, we have presented BLASTER, a framework designed to optimize radio resource management in an integrated TN-NTN. BLASTER aims to control UE association, splits the bandwidth between terrestrial and non-terrestrial tiers, and manages the MBS activation and MBS transmit power level. A novel method for splitting the bandwidth between terrestrial and non-terrestrial tiers is introduced, based on the fraction of UEs associated to the latter. The proposed algorithm also highlights the critical and dynamic part the satellites play in this integrated TN-NTN, adapting their role to various traffic demands. Indeed, the non-terrestrial tier takes a prominent role in low traffic, ensuring that terrestrial MBSs offload their UEs to the satellite to enable their shutdown. Conversely, the non-terrestrial tier embraces a secondary role in high traffic, mainly acting as an outlet to cell-edge UEs, facilitating load distribution while also giving up a share of its resources to the terrestrial tier. Simulation results display notable results, as the average TN energy consumption decreases by $67\,\%$ in low traffic compared to the 3GPP benchmarks, while the average SLT in high traffic increases by roughly $6\,\%$ compared to 3GPP-NTN. In future works, we plan to modify the current deterministic association method, as this method can be challenging to implement in practice. Instead, we aim to develop an ML-based solution, and adopt a distributed approach to tackle the problem. Another potential future enhancement could involve exploring a multi-beam system for the satellites, with a focus on studying inter-beam interference, and how to manage it efficiently, as well as considering the uplink (UL) scenario. Proof of Proposition 1. First, we inject (12) into formula (13) to get:
@@
Indeed, developing the left component of (30), we get:
@@
Then, focusing on the right component of (30), we have: